Gold Mining, Inc. is using the profitability index (PI) when evaluating projects. Gold Mining’s cost of capital is 11.35 percent. What is the PI of a project if the initial costs are $2,118,606 and the project life is estimated as 5 years? The project will produce the same after-tax cash inflows of $479,544 per year at the end of the year.
Answer #1

Profitability Index = Present value of future Cash flows/Initial Investment

Let the cash flow in Year n be CFn

Given,
CF0 = -2118606
CF1 = 479544
CF2 = 479544
CF3 = 479544
CF4 = 479544
CF5 = 479544

Cost of capital = r = 11.35%

Present Value of future cash flows = CF1/(1+r) + CF2/(1+r)2 + CF3/(1+r)3 + CF4/(1+r)4 + CF5/(1+r)5
= 479544/(1+0.1135) + 479544/(1+0.1135)2 + 479544/(1+0.1135)3 + 479544/(1+0.1135)4 + 479544/(1+0.1135)5
= -$1756850.76

Profitability Index = Present value of future Cash flows/Initial Investment

= 1756850.76/2118606

= 0.83